# Nonreciprocal Photon-Phonon Entanglement in Kerr-Modified Spinning Cavity Magnomechanics

### 09 Dec 2023 v1

@@ III.2 Nonreciprocal bistability
From Eq. (III.1), a cubic equation related to the mean magnon number $|m_{s}|^{2}=M$ can be given by
@@
As $K_{0}$ can be tuned via adjusting the direction of the magnetic field, so $K_{0}^{\prime}$ in Eq. (10) can be zero or nonzero. This is because that pure magnomechanical coupling [$m^{\dagger}m(b+b^{\dagger})$], similar to optomechanics, is equivalent to an effective magnon Kerr Hamiltonian by performing the unitary transformation $U=\exp[m^{\dagger}m(b-b^{\dagger})]$ Gong et al. (2008). When $K_{0}^{\prime}=0$, the cubic equation given by Eq. (10) becomes to a linear equation in the mean magnon number $M$. Obviously, it is proportional to the square of the Rabi frequency of the driving field and nonreciprocally responses to the driving fields from opposite directions, as shown in Fig. 2(a). When $K_{0}^{\prime}\neq 0$, the mean magnon number $M$ determined by Eq. (10) can have two switching points for bistability under the specific parameter conditions [see Fig. 2(b)], at which there must be $d\epsilon_{d}/dM=0$, i.e.,
@@
where the positive (negative) simbol denotes $K_{0}^{\prime}>0$ and $\Delta_{m}^{\prime}<0$ ($K_{0}^{\prime}<0$ and $\Delta_{m}^{\prime}>0$). Eq. (14) indicates that magnonic bistability can be predicted when the strength of the driving field exceeds the critical value, i.e., $\epsilon_{d}>\epsilon_{d}^{c}$, as shown in Fig. 2(b). Due to different responses of the spinning cavity to the CW or CCW driving field, nonreciprocal bistability can be apparently observed [see the red and blue curves in Fig. 2(b)].
@@ III.3 Fluctuation dynamics
Apart from the steady state dynamics when the transformation $\sigma\rightarrow\sigma_{s}+\sigma$ is substituted into Eq. (III.1), the fluctation dynamics can also be obtained,
$\dot{a}=$ $-[\kappa_{a}+i(\Delta_{a}-\Delta_{F})]a-ig_{ma}m+\sqrt{2\kappa_{a}}a_{\rm in},$
$\dot{b}=$ $-(\kappa_{b}+i\omega_{b})b-i(G_{mb}m^{\dagger}+G_{mb}^{*}m)+g_{mb}m^{\dagger}m$
@@
$\dot{m}=$ $-[\kappa_{m}+i(\tilde{\Delta}_{m}+2\Delta_{K})]m-ig_{ma}a+\sqrt{2\kappa_{m}}m_{\rm in}$
$-ig_{mb}m(b+b^{\dagger})-2iK_{0}m_{s}^{2}m^{\dagger}-iG_{mb}(b+b^{\dagger})$
$-2iK_{0}(m_{s}^{*}m^{2}+2m_{s}m^{\dagger}m+m^{\dagger}m^{2}),$
where $G_{mb}=g_{mb}m_{s}$ is the effective magnomechanical coupling strength significantly enhanced by multiple magnons. Below we assume that $m_{s}$ is real for simplicity. This can be realized by choosing the propoer phase of the driving field according to Eq. (III.1). Under the strong driving field (i.e., $|m_{s}|\gg 1$), Eq. (III.3) can be linearized by safely ignoring the high-order fluctuations. As a result, Eq. (III.3) reduces to
$\dot{a}=$ $-[\kappa_{a}+i(\Delta_{a}-\Delta_{F})]a-ig_{ma}m+\sqrt{2\kappa_{a}}a_{\rm in},$
$\dot{b}=$ $-(\kappa_{b}+i\omega_{b})b-iG_{mb}(m^{\dagger}+m)+\sqrt{2\kappa_{b}}b_{\rm in},$
$\dot{m}=$ $-[\kappa_{m}+i(\tilde{\Delta}_{m}+2\Delta_{K})]m-i\Delta_{K}m^{\dagger}-ig_{ma}a$
$-iG_{mb}(b+b^{\dagger})+\sqrt{2\kappa_{m}}m_{\rm in}.$ (16)
We then rewrite the above equations as $\dot{\sigma}=i[H_{\rm eff},\sigma]-\kappa_{\sigma}\sigma+\sqrt{2\kappa_{\sigma}}\sigma_{\rm in}$, so the effective Hamiltonian of the linearized system can be given by
$H_{\rm eff}=$ $(\Delta_{a}-\Delta_{F})a^{\dagger}a+\omega_{b}b^{\dagger}b+(\tilde{\Delta}_{m}+2\Delta_{K})m^{\dagger}m$
$+g_{ma}(a^{\dagger}m+am^{\dagger})+\frac{\Delta_{K}}{2}(m^{\dagger}m^{\dagger}+mm)$
$+G_{mb}(m^{\dagger}+m)(b^{\dagger}+b).$ (17)
Note that the two-magnon effect (i.e., $m^{\dagger}m^{\dagger}+mm$) stems from the magnon Kerr nonlinearity in the presence of the strong driving field, which can be well tuned by varying the Rabi frequency of the driving field.
@@ IV The nonreciprocal photon-phonon entanglement
Figure 2: The mean magnon number versus the normalized amplitude of the driving field with (a) $K_{0}=\eta_{b}\omega_{b}$ and (b) $K_{0}=0.1\eta_{b}\omega_{b}$, where $\omega_{b}/2\pi=10$ MHz, $\kappa_{a}=\kappa_{m}=0.1\omega_{b}$, $\kappa_{b}/2\pi=100$ Hz, $g_{\rm ma}=0.2\omega_{b}$, $g_{mb}=10^{-3}\omega_{b}$, $|\Delta_{F}|=0.2\omega_{b}$, $\Delta_{a}=-\omega_{b}$, and $\Delta_{m}=\omega_{b}$.
With the effective Hamiltonian $H_{\rm eff}$ in hand, its dynamics governed by Eq. (III.3) can be rewritten in a more compact form as $\dot{u}(t)=Au(t)+f(t)$, where $u^{T}(t)=(X_{a},Y_{a},X_{m},Y_{m},X_{b},Y_{b})$ is the vector operator of the system, $f^{T}(t)=(\sqrt{2\kappa_{a}}X_{a}^{\rm in}$, $\sqrt{2\kappa_{a}}Y_{a}^{\rm in}$, $\sqrt{2\kappa_{m}}X^{\rm in}_{m}$, $\sqrt{2\kappa_{m}}Y^{\rm in}_{m}$, $\sqrt{2\kappa_{b}}X^{\rm in}_{b}$, $\sqrt{2\kappa_{b}}Y^{\rm in}_{b})$ is the input noise of the system, and
@@
is the drift matrix. Here $X_{\sigma}=(\sigma^{\dagger}+\sigma)/\sqrt{2}$, $Y_{\sigma}=i(\sigma^{\dagger}-\sigma)/\sqrt{2}$, $X_{\sigma}^{\rm in}=(\sigma_{\rm in}^{\dagger}+\sigma_{\rm in})/\sqrt{2}$, and $Y_{\sigma}^{\rm in}=i(\sigma_{\rm in}^{\dagger}-\sigma_{\rm in})/\sqrt{2}$.
@@
Figure 3: Density plot of the photon-phonon entanglement $E_{ab}$ as functions of $\Delta_{a}/\omega_{b}$ and $\tilde{\Delta}_{m}/\omega_{b}$ with (a) $\Delta_{K}>0$, $\Delta_{F}>0$, (b) $\Delta_{K}>0$, $\Delta_{F}<0$, (c) $\Delta_{K}<0$, $\Delta_{F}>0$, and (d) $\Delta_{K}<0$, $\Delta_{F}<0$. Other parameters are the same as those in Fig. 2 except for $G_{mb}=g_{ma}$, $|\Delta_{K}|=0.2\omega_{b}$, and the bath temperature $T=10$ mK.
Figure 4: The photon-phonon entanglement versus $\tilde{\Delta}_{m}/\omega_{b}$ with and without the magnon Kerr effect in the left panel (a, b) and $\Delta_{K}$ and $\Delta_{F}$ in the right panel (c), where (a) $\Delta_{F}>0$ and (b) $\Delta_{F}<0$. Other parameters in (a-c) are the same as those in Fig. 2.
@@
We first plot the logarithmic negativity $E_{ab}$ as functions of the normalized $\Delta_{a}/\omega_{b}$ and $\tilde{\Delta}_{m}/\omega_{b}$ in the presence of both the Sagnac and the magnon Kerr effects in Fig. 3, where $G_{\rm mb}/2\pi=|\Delta_{k}|=0.2\omega_{b}$, the bath temperature $T=10$ mK, and other parameters are the same as those in Fig. 2. The chosen parameters ensure the system stable following the Routh-Hurwitz criterion. From Fig. 3, one can see that the photon-phonon entanglement can be tuned by changing the frequency detunings $\Delta_{a}$ and $\tilde{\Delta}_{m}$. In particular, its optimal value is predicted at $\Delta_{a}\approx-\omega_{b}+\Delta_{F}$ and $\tilde{\Delta}_{m}\approx\omega_{b}-2\Delta_{K}$. The mechanism of this optimal entanglement can be interpreted as follows: When $\tilde{\Delta}_{m}\approx\omega_{b}-2\Delta_{K}$, the magnomechanical subsystem is driven to the red sideband, where the mechanical mode can be well cooled for allowing considerable magnon-phonon entanglement owing to the enhanced magnomechanical coupling $G_{mb}$. Then the magnon-phonon entanglement is significantly transferred to the photon-phonon entanglement via the beam splitter magnon-photon interaction at $\Delta_{a}\approx-\omega_{b}+\Delta_{F}$. From Fig. 3, we find that the predicted photon-phonon entanglement nonreciprocally response to the change of the frequency detunings $\Delta_{K}$ or $\Delta_{F}$, i.e., the magnon Kerr or the Sagnac effect. This means that the nonreciprocal photon-photon entanglement can be achieved by including these two effects. Specifically, when $\Delta_{F}>0$, the optimal cavity frequency detuning is fixed at $\Delta_{a}\approx-0.8\omega_{b}$, but the optimal magnon frequency detuning is $\tilde{\Delta}_{m}\approx 0.6\omega_{b}$ for $\Delta_{K}>0$ [Fig. 3(a)] and $\tilde{\Delta}_{m}\approx 1.4\omega_{b}$ for $\Delta_{K}<0$ [Fig. 3(c)]. This indicates that the photon-phonon entanglement nonreciprocally responses to the opposite magnetic fields when the magnon Kerr effect is considered with the fixed Sagnac effect. The similar result can also be obtained from Figs. 3(b) and 3(d). By comparing Figs. 3(a) with 3(b) [or Figs. 3(c) with 3(d)], we show that the optimal photon-phonon entanglement is predicted at different cavity frequency detunnings $\Delta_{a}$ but at the same magnon frequency detuning $\tilde{\Delta}_{m}$. This means the nonreciprocal photon-phonon entanglement can also be induced by the Sagnac effect for the fixed magnetic field in the presence of the magnon Kerr effect. Figures 4(a) and 4(b) futher show the nonreciprocal behavior of the photon-phonon entanglement with the normalized $\tilde{\Delta}_{m}/\omega_{b}$ in the presence of the magnon and Sagnac effects, where $\Delta_{a}=-\omega_{b}$. For $\Delta_{F}>0$ [see Fig. 4(a)], we find that the photon-phonon entanglement can be nonreciprocally enhanced ($\Delta_{K}>0$) or reduced ($\Delta_{K}>0$), compared to the case without the magnon Kerr effect. In the case of $\Delta_{F}<0$ [see Fig. 4(b)], the situation becomes opposite. From Fig. 4(a) and 4(b), one can see that the photon-phonon entanglement can also be nonreciprocally enhanced or reduced by the Sagnac effect for the fixed magnon Kerr effect. Figure 4(c) directly reveals the behavior of the photon-phonon entanglement with these two effects. Obviously, the nonreciprocal photon-phonon entanglement can be achieved by using the magnon Kerr effect or the Sagnac effect or both of them. The optimal value of the photon-phonon entanglement can be approximately predicted at $\Delta_{K}\Delta_{F}<0$.
@@
Figure 5: The photon-phonon entanglement versus $G_{mb}/g_{ma}$ with (a) $\Delta_{K}=0$, (b) $\Delta_{K}>0$, (c) $\Delta_{K}<0$, wehre the Sagnac effect are all considered. (d) The nonreciprocity of the photon-phonon entanglement induced by the magnon Kerr effect versus $G_{mb}/g_{ma}$ with and without the Sagnac effect. Other parameters in (a-c) are the same as those in Fig. 2.

### 12 Apr 2024 v2

@@ III.2 Nonreciprocal bistability
From Eqs. (III.1), a cubic equation related to the mean magnon number $|m_{s}|^{2}=M$ can be given by
@@
Figure 2: The mean magnon number vs the normalized amplitude of the driving field with (a) $K_{0}=\eta_{b}\omega_{b}$ and (b) $K_{0}=0.1\eta_{b}\omega_{b}$, where $\omega_{b}/2\pi=10$ MHz, $\kappa_{a}=\kappa_{m}=0.1\omega_{b}$, $\kappa_{b}/2\pi=100$ Hz, $g_{\rm ma}=0.2\omega_{b}$, $g_{mb}=10^{-3}\omega_{b}$, $|\Delta_{F}|=0.2\omega_{b}$, $\Delta_{a}=-\omega_{b}$, and $\Delta_{m}=\omega_{b}$.
As $K_{0}$ can be tuned via adjusting the direction of the magnetic field, so $K_{0}^{\prime}$ in Eq. (10) can be zero or nonzero. This is because pure magnomechanical coupling [$m^{\dagger}m(b+b^{\dagger})$], similar to optomechanics, is equivalent to an effective magnon Kerr Hamiltonian by performing the unitary transformation $U=\exp[m^{\dagger}m(b-b^{\dagger})]$ Gong et al. (2009). When $K_{0}^{\prime}=0$, the cubic equation given by Eq. (10) becomes a linear equation in the mean magnon number $M$. Obviously, it is proportional to the square of the Rabi frequency of the driving field and nonreciprocally responses to the driving fields from opposite directions, as shown in Fig. 2(a). When $K_{0}^{\prime}\neq 0$, the mean magnon number $M$ determined by Eq. (10) can have two switching points for bistability under specific parameter conditions [see Fig. 2(b)], at which there must be $d\epsilon_{d}/dM=0$, i.e.,
@@
where the positive (negative) symbol denotes $K_{0}^{\prime}>0$ and $\Delta_{m}^{\prime}<0$ ($K_{0}^{\prime}<0$ and $\Delta_{m}^{\prime}>0$). Eq. (14) indicates that magnonic bistability can be predicted when the strength of the driving field exceeds the critical value, i.e., $\epsilon_{d}>\epsilon_{d}^{c}$, as shown in Fig. 2(b). Due to different responses of the spinning cavity to the CW or CCW driving field, nonreciprocal bistability can be apparently observed [see the red and blue curves in Fig. 2(b)].
@@ III.3 Fluctuation dynamics
Apart from the steady-state dynamics when the transformation $\sigma\rightarrow\sigma_{s}+\delta\sigma$ is substituted into Eqs. (III.1), the fluctuation dynamics can also be obtained:
$\delta\dot{a}=$ $-[\kappa_{a}+i(\Delta_{a}-\Delta_{F})]\delta a-ig_{ma}\delta m+\sqrt{2\kappa_{a}}a_{\rm in},$
$\delta\dot{b}=$ $-(\kappa_{b}+i\omega_{b})\delta b-i(G_{mb}\delta m^{\dagger}+G_{mb}^{*}\delta m)+g_{mb}\delta m^{\dagger}\delta m$
@@
$\delta\dot{m}=$ $-[\kappa_{m}+i(\tilde{\Delta}_{m}+2\Delta_{K})]\delta m-ig_{ma}\delta a+\sqrt{2\kappa_{m}}m_{\rm in}$
$-ig_{mb}\delta m(\delta b+\delta b^{\dagger})-2iK_{0}m_{s}^{2}\delta m^{\dagger}-iG_{mb}(\delta b+\delta b^{\dagger})$
$-2iK_{0}(m_{s}^{*}\delta m^{2}+2m_{s}\delta m^{\dagger}\delta m+\delta m^{\dagger}\delta m^{2}),$
where $G_{mb}=g_{mb}m_{s}$ is the effective magnomechanical coupling strength significantly enhanced by multiple magnons. Below we assume that $m_{s}$ is real for simplicity. This can be realized by choosing the proper phase of the driving field according to Eqs. (III.1). Under the strong driving field, the condition $|a_{s}|\gg 1$ can be realized. Because of the beam-splitter interaction ($\propto a^{\dagger}m+am^{\dagger}$) between the photons and magnons, we have $|m_{s}|\gg 1$, which directly requires the system parameters to satisfy
$\varepsilon_{d}^{2}g_{ma}^{2}\gg|[\kappa_{m}+i(\tilde{\Delta}_{m}+\Delta_{k})][\kappa_{a}+i(\Delta_{a}-\Delta_{F})]+g_{ma}^{2}|^{2}$ (16)
according to Eqs. (III.1). This indicates that the mean-field approximation can be applied to Eqs. (III.3) when Eq. (16) is kept, so that the high-order fluctuations in Eq. (III.3) can be safely ignored. As a result, Eqs. (III.3) reduces to
$\delta\dot{a}=$ $-[\kappa_{a}+i(\Delta_{a}-\Delta_{F})]\delta a-ig_{ma}\delta m+\sqrt{2\kappa_{a}}\delta a_{\rm in},$
$\delta\dot{b}=$ $-(\kappa_{b}+i\omega_{b})\delta b-iG_{mb}(\delta m^{\dagger}+\delta m)+\sqrt{2\kappa_{b}}b_{\rm in},$
$\delta\dot{m}=$ $-[\kappa_{m}+i(\tilde{\Delta}_{m}+2\Delta_{K})]\delta m-i\Delta_{K}\delta m^{\dagger}-ig_{ma}\delta a$
$-iG_{mb}(\delta b+\delta b^{\dagger})+\sqrt{2\kappa_{m}}m_{\rm in}.$ (17)
We then rewrite the above equations as $\delta\dot{\sigma}=i[H_{\rm eff},\delta\sigma]-\kappa_{\sigma}\delta\sigma+\sqrt{2\kappa_{\sigma}}\sigma_{\rm in}$, so the effective Hamiltonian of the linearized system can be given by
$H_{\rm eff}=$ $(\Delta_{a}-\Delta_{F})\delta a^{\dagger}\delta a+\omega_{b}\delta b^{\dagger}b+(\tilde{\Delta}_{m}+2\Delta_{K})\delta m^{\dagger}\delta m$
$+g_{ma}(\delta a^{\dagger}\delta m+\delta a\delta m^{\dagger})+\frac{\Delta_{K}}{2}(\delta m^{\dagger}\delta m^{\dagger}+\delta m\delta m)$
$+G_{mb}(\delta m^{\dagger}+\delta m)(\delta b^{\dagger}+\delta b).$ (18)
Note that the two-magnon effect (i.e., $\delta m^{\dagger}\delta m^{\dagger}+\delta m\delta m$) stems from the magnon Kerr nonlinearity in the presence of the strong driving field, which can be well tuned by varying the strength of the driving field.
@@ IV The nonreciprocal photon-phonon entanglement
IV.1 Entanglement calculation
With the effective Hamiltonian $H_{\rm eff}$ in hand, its dynamics governed by Eqs. (III.3) can be rewritten in a more compact form as $\dot{u}(t)=Au(t)+f(t)$, where $u^{T}(t)=(X_{a},Y_{a},X_{m},Y_{m},X_{b},Y_{b})$ is the vector operator of the system, $f^{T}(t)=(\sqrt{2\kappa_{a}}X_{a}^{\rm in}$, $\sqrt{2\kappa_{a}}Y_{a}^{\rm in}$, $\sqrt{2\kappa_{m}}X^{\rm in}_{m}$, $\sqrt{2\kappa_{m}}Y^{\rm in}_{m}$, $\sqrt{2\kappa_{b}}X^{\rm in}_{b}$, $\sqrt{2\kappa_{b}}Y^{\rm in}_{b})$ is the input noise of the system, and
@@ IV.1 Entanglement calculation
is the drift matrix. Here $X_{\sigma}=(\delta\sigma^{\dagger}+\delta\sigma)/\sqrt{2}$, $Y_{\sigma}=i(\delta\sigma^{\dagger}-\delta\sigma)/\sqrt{2}$, $X_{\sigma}^{\rm in}=(\sigma_{\rm in}^{\dagger}+\sigma_{\rm in})/\sqrt{2}$, and $Y_{\sigma}^{\rm in}=i(\sigma_{\rm in}^{\dagger}-\sigma_{\rm in})/\sqrt{2}$.
@@
IV.2 Nonreciprocal entanglement exploration
We first plot the logarithmic negativity $E_{ab}$ as functions of the normalized $\Delta_{a}/\omega_{b}$ and $\tilde{\Delta}_{m}/\omega_{b}$ in the presence of both the Sagnac and the magnon Kerr effects in Fig. 3. The chosen parameters are the same as those in Fig. 2 except for $G_{mb}=0.2\omega_{b}$, $|\Delta_{F}|=|\Delta_{K}|=0.1\omega_{b}$ and the bath temperature $T=10$ mK. These parameters can ensure the system is stable according to the Routh-Hurwitz criterion.
@@ IV.2 Nonreciprocal entanglement exploration
Figure 3: Density plot of the photon-phonon entanglement $E_{ab}$ as functions of $\Delta_{a}/\omega_{b}$ and $\tilde{\Delta}_{m}/\omega_{b}$ with (a) $\Delta_{K}>0$, $\Delta_{F}>0$, (b) $\Delta_{K}>0$, $\Delta_{F}<0$, (c) $\Delta_{K}<0$, $\Delta_{F}>0$, and (d) $\Delta_{K}<0$, $\Delta_{F}<0$. Other parameters are the same as those in Fig. 2 except for $G_{mb}=0.2\omega_{b}$, $|\Delta_{F}|=|\Delta_{K}|=0.1\omega_{b}$, and the bath temperature $T=10$ mK.
@@
From Fig. 3, one can see that the photon-phonon entanglement can be tuned by changing the frequency detunings $\Delta_{a}$ and $\tilde{\Delta}_{m}$. In particular, its optimal value is predicted at $\Delta_{a}\approx-\omega_{b}+\Delta_{F}$ and $\tilde{\Delta}_{m}\approx\omega_{b}-2\Delta_{K}$. The mechanism of this optimal entanglement can be interpreted as follows: When $\tilde{\Delta}_{m}\approx\omega_{b}-2\Delta_{K}$, the magnomechanical subsystem is driven to the red sideband, where the mechanical mode can be well cooled for allowing considerable magnon-phonon entanglement owing to the enhanced magnomechanical coupling $G_{mb}$. Then the magnon-phonon entanglement is significantly transferred to the photon-phonon entanglement via the beam-splitter magnon-photon interaction at $\Delta_{a}\approx-\omega_{b}+\Delta_{F}$. Moreover, from Fig. 3, we also find that the predicted photon-phonon entanglement nonreciprocally responds to the change of the frequency detunings $\Delta_{K}$ or $\Delta_{F}$, i.e., the magnon Kerr or the Sagnac effects. This means that the nonreciprocal photon-photon entanglement can be achieved by including these two effects. Specifically, when $\Delta_{F}>0$, the optimal cavity frequency detuning is fixed at $\Delta_{a}\approx-0.8\omega_{b}$, but the optimal magnon frequency detuning is $\tilde{\Delta}_{m}\approx 0.6\omega_{b}$ for $\Delta_{K}>0$ [Fig. 3(a)] and $\tilde{\Delta}_{m}\approx 1.4\omega_{b}$ for $\Delta_{K}<0$ [Fig. 3(c)]. This indicates that the photon-phonon entanglement nonreciprocally changes with the magnon Kerr effect when the Sagnac effect is fixed. A similar result can also be obtained from Figs. 3(b) and 3(d). By comparing Figs. 3(a) with 3(b) [or Figs. 3(c) with 3(d)], the optimal photon-phonon entanglement has left shifts on the frequency detuning $\Delta_{a}$, which means the nonreciprocal photon-phonon entanglement can be induced by the Sagnac effect when the magnon Kerr effect is fixed.
@@
Figure 4: The photon-phonon entanglement vs $\tilde{\Delta}_{m}/\omega_{b}$ with and without the magnon Kerr effect (a, b) and $\Delta_{K}$ and $\Delta_{F}$ (c), where (a) $\Delta_{F}>0$ and (b) $\Delta_{F}<0$. Other parameters in panels (a)-(c) are the same as those in Fig. 3.
Figures 4(a) and 4(b) further show the nonreciprocal behavior of the photon-phonon entanglement with the normalized $\tilde{\Delta}_{m}/\omega_{b}$ with both the nonlinear effects, where $\Delta_{a}=-\omega_{b}$ is fixed. For $\Delta_{F}>0$ [see Fig. 4(a)], we find that the photon-phonon entanglement can be nonreciprocally enhanced ($\Delta_{K}>0$) or reduced ($\Delta_{K}>0$), compared to the case without the magnon Kerr effect. In the case of $\Delta_{F}<0$ [see Fig. 4(b)], the situation becomes opposite. From Fig. 4(a) and 4(b), one can see that the photon-phonon entanglement can also be nonreciprocally enhanced or reduced by the Sagnac effect with the magnon Kerr effect. Figure 4(c) directly shows the behavior of the photon-phonon entanglement with these two effects. We show that the nonreciprocal photon-phonon entanglement can be predicted in a broad range of $\Delta_{K}$ and $\Delta_{F}$, and its optimal value can be obtained around the region of $\Delta_{K}\Delta_{F}<0$.
IV.3 The magnomechanical coupling effect